OpenStudy (anonymous):

A cylindrical cork is 6 cm high and has a TSA of 33.38 cm2. Could it be used to close a bottle whose neck is 3 cm in diameter? Justify your answer.

OpenStudy (anonymous):

sorry but that is the incorrect answer.

OpenStudy (anonymous):

TSA = 2pi*r^2 + 2pi*r*h 33.38 = 6.28r^2 + (6.38)(6)*r 33.38 = 6.28r^2 + 37.7r 0 = 6.28r^2 + 37.7r - 33.38 i believe you use quadratic formula and take the positive root

OpenStudy (anonymous):

the radius of the bottle is 3/2 = 1.5 so compare that with the radius of the cork

OpenStudy (anonymous):

the answers say that the diameter of the cork is 1.6cm though.

OpenStudy (anonymous):

I want to know how 1.6cm is achieved.

OpenStudy (anonymous):

and the radius of the bottle is indeed 1.5 but we don't need that, we are comparing the diameters.

OpenStudy (anonymous):

you can compare diameters but you can also compare radii

OpenStudy (anonymous):

since TSA is in terms of radius, i just turned diameter into radius

OpenStudy (anonymous):

yep but how is 1.6cm of the diameter of the cork obtained?

OpenStudy (anonymous):

like i said use quadratic formula on 6.28r^2 + 37.7r - 33.38 = 0 have you tried that?

OpenStudy (anonymous):

tell me when you arrive with an answer. remember to take the positive root

OpenStudy (anonymous):

0.78?

OpenStudy (anonymous):

k, I get it now, thanks.

OpenStudy (anonymous):

you're welcome